OpenStudy (anonymous):

Can someone teach me to do these? y varies directly as the square of x. When x = 4, y = 8. Find x when y = 32. x = ±4 x = ±6 x = ±8 x = ±12

OpenStudy (anonymous):

X=kY find K by first 2 values and then put it in the above eq for finding the remaining values of Y

OpenStudy (anonymous):

what is k ?

OpenStudy (anonymous):

the proportionality constant K which u put after removing the sign of proportionality

zepdrix (zepdrix):

y varies directly as the `square` of x.\[\Large y=k x^2\] I'm confused by your equation hamza :o

OpenStudy (anonymous):

y=kx^2 when x=4, y=8 8=k*4^2=16k k=8/16=1/2

OpenStudy (anonymous):

o sorry i had missed the square

OpenStudy (anonymous):

when y=32 32=1/2 x^2 x^2=64 \[x=\pm 8\]

OpenStudy (jdoe0001):

when something varies directly to something else, means if a = 5, b = 10 if a changes to 10, b changes to 20 if a changes to 20, b changes to 40 and so on "varies directly" means, is being affected directly by the change on some other fellow as "a" changes, the changes ripple to "b" but if "a" change say 10 units, it doesn't have to ripple or reflect on "b" to another extra 10 units, it changes with a multiplier value, let's call it "k" b = (k) a if a =5 then b = (k) (5) then b changes to 10, so what's the multiplier "k" 10 = (k) (5) #solving for "k" 10/5 = k 2 = k if a = 10, then b is b = (2)10 b = 20 and so on

zepdrix (zepdrix):

The k can be a little confusing if you don't understand the concept. Like if I told you that `money` is directly related to `pizza`, I would represent it this way:\[\large M=kP\]Where \(\large k\) is maybe the cost of each slice of pizza. Oh man.. that was the worst example ever lol.. ignore this XD

OpenStudy (jhannybean):

\[\large y = kx^2\]\[\large 8= (4)^2k\]\[\large 8 = 16k\]\[\large \frac{1}{2} = k\]\[\large y= \frac{x^2}{2}\]\[\large 32 = \frac{x^2}{2}\]\[\large 32 \cdot 2 =x^2\]\[\large \sqrt{64} = x\]\[\large x =\ ? \ \]

OpenStudy (anonymous):

@Jhannybean can u tell me how to type like this as u did?

OpenStudy (jhannybean):

Equation editor. Follow @jdoe0001 's explanation, then look back at mine. It'll make moresense.

OpenStudy (jdoe0001):

@hamza10 http://www.onemathematicalcat.org/MathJaxDocumentation/TeXSyntax.htm you can try and test at => http://mathb.in/

OpenStudy (anonymous):

Woah thats alot to understand lol. @Jhannybean will do .

OpenStudy (anonymous):

so x = 8?

zepdrix (zepdrix):

yes, \(\pm8\) :)

OpenStudy (anonymous):

thank you :)